OpenStudy (anonymous):

(please help :( I'll give a fan and a medal ) 1. Write a function rule to represent the following: Two less than y is 8 more than the product of 3 and x a. 2y= 8x + 3 b. y + 2= 3x + 8 c. y - 2 = 3x + 8 d. y- 2 = 8 - 3x

OpenStudy (cj49):

two less than y=y-2 is means equal to 8 more means 8 plus the prodouct means multiplication.

OpenStudy (anonymous):

:o ohh...wait what? I'm lost again

OpenStudy (cj49):

what u din understood mate...

OpenStudy (anonymous):

i think its c

OpenStudy (cj49):

pls dun provide direct ans@torijordan113

OpenStudy (anonymous):

I mean, C is what I was thinking. I just didn't know if I was right or not :/

OpenStudy (cj49):

k

OpenStudy (cj49):

its C bc two less than y=y-2 is 8 more than the product of 3x is=8+3x

OpenStudy (anonymous):

ohh, oki. thankies. :). May I ask some more please?

OpenStudy (cj49):

sure mate

OpenStudy (anonymous):

thank you! okay 2. Write a function rule that gives the total cost c(p) of p pounds of sugar if each pround cost $0.59 a. c(p)= 59p b. c(p)= p\[\frac{ p }{ 59 }\] c. c(p)= p + 0.59 d. c(p) = 0.59

OpenStudy (cj49):

.59x total pounds purchased =.59xp

OpenStudy (anonymous):

how did you get that? :o

OpenStudy (cj49):

we would get total cost by multplying the total no on quantity to the amt of per pound let P be the pounds of sugar and each pound costs .59 therefore total cpst=.59xP eg if one pencil would cost $1 then how much would 5 pencils cost = 5x1 =total cost $5

OpenStudy (anonymous):

ohhh! oki! I see :)

OpenStudy (anonymous):

so the answer will be A :D?

OpenStudy (cj49):

yups

OpenStudy (anonymous):

yayyyy :D

OpenStudy (anonymous):

gosh, you're so sweet and amazing ♥♥

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

<3 <3 <3 <3 <3 <3 \(^0^)/